nekono12:

i. To see if the slopes are the same, write an equation setting the two fractions equal to each other. Is this equation true? Why or why not? What does that mean about the slope between points E and A and the slope between the points A and C? HELP ME PLEEASE!

nekono12:

heres the photo if needed!

Luigi0210:

If you're looking for the slopes of AC and AE, you're gonna need to use the following points: A(0,0) C(3, 2) A(0,0) E(-6,-4) \[\Large m_{AC} = \frac{2-0}{3-0}=\frac{2}{3} \] \[\Large m_{AE}=\frac{-4-0}{-6-0}=\frac{-4}{-6}=\frac{2}{3}\]

nekono12:

BRO THANK YOU

Luigi0210:

You're welcomee, as you can see, they're the same slopes.